​(Weighted average cost of capital​) Crawford Enterprises is a publicly held company located in​ Arnold, Kansas. The firm began as a small tool and die shop but grew over its​ 35-year life to become a leading supplier of metal fabrication equipment used in the farm tractor industry. At the close of​ 2019, the​ firm's balance sheet appeared as​ follows:

Cash   460,000      
Accounts receivable   3,910,000      
Inventories   8,200,000   Long-term debt   11,270,000
Net property, plant, and equipment   17,715,000   Common equity   19,015,000
Total assets   30,285,000   Total debt and equity   30,285,000

At present the​ firm's common stock is selling for a price equal to its book​ value, and the​ firm's bonds are selling at par.​ Crawford's managers estimate that the market requires a return of 17 percent on its common​ stock, the​ firm's bonds command a yield to maturity of 8 ​percent, and the firm faces a tax rate of 28 percent.

a. What is​ Crawford's weighted average cost of​ capital?

b. If​ Crawford's stock price were to rise such that it sold at 1.5 times book​ value, causing the cost of equity to fall to 15 ​percent, what would the​ firm's cost of capital be​ (assuming the cost of debt and tax rate do not​ change)?

c. Crawford is considering a new business opportunity involving the acquisition of a trucking firm. What do you think the firm should do to select an appropriate cost of capital for evaluating this​ acquisition?
Answer #1

a. As per the given information,

Pre-tax Cost of Debt = 8% ( bonds are trading at YTM of 8%)

Cost of Equity = 17% (return on common equity is 17% which if the company plans to purchase same equity will have to pay)

Debt to Enterpise value (D/V) = 11270000/(30285000-460000). ( Enterpise value = Equity + Debt -Cash)

= 11270000/29825000 = 37.8%

Weighted Average Cost of Capital = 17%*(1-37.8%) + 8%*(1-28%)*37.8%

= 12.75%

b. We should note that though the stock is selling at 1.5 times book value, the value is not realised. Hence, total debt and equity on the books still remain the same.

Based on the above, the only change to be done to the formula in a is cost of equity to 15%.

WACC = 15%*(1-37.8%) + 8%*(1-28%)*37.8%

= 11.5%

c. Management should calculate the expected weighted rate of capital of the trucking firm based on the trucking firm's capital costs as well as what other market competitors are willing to pay for the trucking firm. Based on the value returned in both cases, Management should be able to select an appropriate cost of capital.